This is the beginning of an article on L^{A}TEX! There are bold text, italics, \underline{underscoring}, footnotes^{1} - and lots of other goodies. A math example:

∑_{n=1}^{∞} \frac{1}{n^{2}} = \frac{π^{2}}{6} .

1this is a footnote